OpenStudy (anonymous):

Which of the following tables shows the correct steps to transform x2 + 4x + 3 = 0 into the form (x - p)2 = q?

OpenStudy (anonymous):

cannot submit the possible answers sorry ill try actually

OpenStudy (anonymous):

Step 1 x2 + 4x + 3 + 1 = 0 + 1 Step 2 x2 + 4x + 4 = 1 Step 3 (x + 2)2 = 1 Step 1 x2 + 4x + 3 - 1 = 0 - 1 Step 2 x2 + 4x + 2 = -1 Step 3 (x + 2)2 = -1 Step 1 x2 + 4x + 3 - 2 = 0 - 2 Step 2 x2 + 4x + 1 = -2 Step 3 (x + 2)2 = -2 Step 1 x2 + 4x + 3 + 2 = 0 + 2 Step 2 x2 + 4x + 5 = 2 Step 3 (x + 2)2 = 2

OpenStudy (anonymous):

@ganeshie8 @Hero @iambatman

OpenStudy (anonymous):

@dumbcow @thomaster @phi @amistre64

OpenStudy (phi):

can you multiply (x+2)(x+2) ? what do you get ?

OpenStudy (anonymous):

x^2+4x+4

OpenStudy (phi):

that means you should look for the choice that replaces x^2+4x+4 with (x+2)^2

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So A?

OpenStudy (phi):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

:3 algebra + 6th grade is a bad mix xD

OpenStudy (anonymous):

no matter how brainy you are :3